### v2

# Analysis and Optimized CXL-Attached Memory Allocation for Long-Context LLM Fine-Tuning

Abstract
The substantial memory requirements of Large Language Models (LLMs), particularly for long-context fine-tuning, have renewed interest in CPU offloading to augment limited GPU memory. However, as context lengths grow, relying on CPU memory for intermediate states introduces a significant bottleneck that can exhaust the capacity of mainstream client platforms. To address this limitation, this work investigates the effectiveness of Compute Express Link (CXL) add-in card (AIC) memory as an extension to CPU memory, enabling larger model sizes and longer context lengths during fine-tuning. Extensive benchmarking reveals two critical challenges. First, current deep learning frameworks such as PyTorch lack fine-grained, per-tensor control over NUMA memory allocation, exposing only coarse, process-level policies. Second, due to this lack of control, when the memory footprint of fine-tuning is offloaded across local DRAM and CXL-attached memory, naively placing optimizer data in higher-latency CXL leads to substantial slowdowns in the optimizer step (e.g., $\sim$4× once data exceeds $\sim$20M elements). To overcome these challenges, this work introduces a PyTorch extension that enables tensor-level system memory control and a CXL-aware memory allocator that pins latency‑critical tensors in local DRAM while maximizing bandwidth by striping latency‑tolerant tensors across one or more CXL devices. Evaluated on a real hardware setup with 7B and 12B models, 4K–32K contexts, and a single GPU, our approach recovers throughput to 97–99% of DRAM‑only with a single AIC and $\approx$100% with two AICs, delivering up to 21% improvement over naive interleaving while preserving DRAM‑like DMA bandwidth for GPU transfers. These results show that carefully managed CXL‑attached memory is a practical path to scaling long‑context fine‑tuning beyond DRAM limits.
I Introduction
The rapid growth of Large Language Models (LLMs) and their ever-increasing parameter counts have introduced significant challenges in memory capacity [1]. As these models frequently exceed available GPU memory, performance bottlenecks arise during both training and deployment. Memory requirements are further exacerbated by the push toward longer context lengths [2], which arise from applications such as long chain-of-thought reasoning [3, 4], generative agents [5], in-context learning [6], retrieval-augmented generation [7], and multimodal tasks [8], all of which are experiencing rapid growth. To support these scenarios, fine-tuning LLMs on long-context datasets has become increasingly important [9, 10, 11, 12, 13]. However, long-context fine-tuning imposes substantial memory overhead, primarily from storing intermediate activation values that scale with context length [14, 15]. Furthermore, limited memory resources restrict batch size during training, thereby constraining throughput.
To address these constraints, particularly in resource-limited environments, offloading strategies, such as CPU offloading and solid-state drive (SSD) offloading, have been proposed [16, 17]. CPU offloading migrates model states, such as parameters, gradients, optimizer data, and occasionally checkpointed activations, from GPU memory to system memory (see Figure 1), while SSD offloading further offloads these states onto SSDs, leveraging the larger and more cost-effective capacity of SSDs. However, SSD offloading suffers from inherent performance and endurance limitations of NAND flash memory, making CPU offloading the more practical and widely adopted approach [16, 18, 19, 20]. Although CPU offloading enables the fine-tuning of larger models and supports longer context lengths on GPU-memory-constrained systems, system memory capacity itself become the bottleneck instead, especially as model sizes continue to grow and the demand for longer contexts and larger batch sizes increases.
System memory capacity on mainstream client platforms (often  192–256 GB today [21]) is constrained by CPU/chipset limits, DIMM slots, and module density. To overcome these constraints, Compute Express Link (CXL) technology has emerged as a promising alternative, providing a viable solution to memory bottlenecks encountered during CPU offloaded long-context LLM fine-tuning [22, 23, 24]. Leveraging PCIe and DRAM, CXL-attached memory expands host capacity beyond DIMM density/slot limits without the performance penalties of NAND-flash SSDs or the cost of high-capacity DIMMs [25]. In particular, CXL Type-3 add-in cards (AICs) provide memory expansion [26, 25] that the host OS typically exposes as CPU-less NUMA nodes, allowing applications to access them similarly to remote DRAM, though with distinct performance characteristics. This capacity enables long-context LLM fine-tuning without being limited by the system memory size; however, since CPU offloading workloads are sensitive to system memory access latency, naively integrating CXL-attached memory into existing CPU offloading workflows does not inherently ensure optimal performance.
Consequently, custom CXL memory management policies tailored to offloading workloads are required [27, 28]. Recent work has begun exploring CXL-attached memory for LLM workloads. For example, Wang et al. [24] evaluate end-to-end performance of CPU offloading with CXL, while Tang et al. [29] employ CXL memory to store the KV cache during inference. However, these studies mainly characterize general performance without analyzing workload-specific behavior or proposing optimizations tailored to CPU offloaded long-context LLM fine-tuning tasks. This leaves an open gap in understanding interactions such as frequent GPU–CPU transfers and latency-sensitive optimizer phases during long-context fine-tuning.
In practice, CPU offloading techniques such as ZeRO-Offload [16] reduce GPU memory consumption by transferring model parameters, gradients, and optimizer states to system memory. While this effectively alleviates GPU pressure, it introduces frequent data transfers between GPU and CPU, and exposes performance sensitivity to memory access latency. When CXL-attached memory is used as an extension or replacement for local DRAM in offloading scenarios, its distinct performance characteristics, including higher latency and lower bandwidth compared to local DIMMs, need to be carefully considered [30]. Generic operating system (OS)-level mechanisms, such as tiered memory systems [31, 32] or interleaving policies [33], provide transparent support but often yield suboptimal performance for specialized workloads. To clarify these implications, this work benchmarks CXL-attached memory under long-context CPU offloading workloads and identifies two primary performance challenges and one key performance characteristic.
First, current deep learning frameworks such as PyTorch [34] enforce a single uniform allocation policy across DRAM and CXL-attached memory. Because memory is managed only at the process level, users are limited to coarse-grained tools like numactl [35], making fine-grained, per-tensor placement impossible and often leading to latency-sensitive tensors being allocated in CXL-attached memory. Second, CPU-based optimizer phases are particularly latency sensitive: read–modify–write loops over parameters, gradients, and optimizer states degrade significantly if data reside in CXL-attached memory rather than DRAM, as the memory accesses of the CPU-based optimizer step dominate the critical path and directly reduce throughput. On the other hand, this work identifies a key performance characteristic: in the GPU pipeline, asynchronous DMA overlaps data movement with kernel execution [^1] A kernel is a low-level function that performs a specific tensor operation (e.g., convolution, matrix multiply). Kernel execution is running that function on hardware (CPU/GPU) to carry out the computation in parallel., and because both DRAM and CXL-attached memory traverse the same PCIe path, host-to-device transfer throughput is broadly comparable from either source.
To address these challenges and build on insights from prior analysis, this work introduces two primary optimizations. First, a fine-grained memory controller is designed and implemented as a PyTorch extension to enable per-tensor control over NUMA memory placement. Second, this paper introduces a CXL-aware memory allocator, implemented as a greedy policy that leverages the fine-grained memory controller to partition tensors by latency sensitivity. Latency-critical data are placed in local DRAM, while latency-tolerant data are directed to CXL-attached memory, with interleaving applied when beneficial to maximize bandwidth. In other words, through the introduced components, latency-bound optimizer states/updates in CPU-offloaded fine-tuning are pinned in DRAM, while bandwidth-bound GPU transfers are striped over CXL AICs via tensor-level placement. Together, these optimizations demonstrate that CXL-attached memory can effectively expand capacity for long-context LLM fine-tuning while achieving performance nearly identical to DRAM-only configurations. Across 7B/12B models and 4K–32K contexts with a single GPU, our CXL‑aware memory allocator restores single‑AIC throughput to 97–99% of DRAM‑only and matches DRAM‑only with dual AICs, outperforming naive interleaving by up to 21%. The main contributions of this study are as follows.
  1. 1.
    To the best of our knowledge, this paper presents the first empirical characterization of CXL-attached memory for long-context LLM fine-tuning, identifying and analyzing the key performance bottlenecks (e.g., optimizer latency sensitivity) introduced by naive CXL adoption.
  2. 2.
    A PyTorch extension is implemented to enable precise, per-tensor memory placement on specified NUMA nodes, a crucial capability for heterogeneous memory systems.
  3. 3.
    A CXL-aware memory allocator is introduced as a greedy policy that maps tensors by latency sensitivity and leverages interleaving when beneficial to maximize bandwidth and minimize latency-induced slowdowns.
  4. 4.
    Real-world experimental results on CXL devices demonstrate that CXL-attached memory can expand capacity while delivering performance comparable to DRAM-only baselines for long-context LLM fine-tuning workloads.
The remainder of the paper is structured as follows: Section II provides background and related work; Section III analyzes CXL-attached memory performance; Section IV details our proposed optimizations; Section V presents experimental evaluations; Section VI reports related works and Section VII concludes this study.
II Background
This section provides background on CPU offloading techniques for long-context LLM fine-tuning (See Section II-A), highlighting the associated system memory bottlenecks (See Section II-B) and the role of Compute Express Link (CXL) technology in addressing these limitations (See Section II-C).
II-A CPU Offloading for Long-Context Fine-Tuning
ZeRO-Offload [16] is a widely used technique to train LLMs on systems with limited GPU memory. It conserves GPU resources by transferring model parameters, gradients, and optimizer states from GPU memory to system memory, and only retrieves them back to GPU memory when required for computation. To further reduce memory usage, ZeRO-Offload can be combined with techniques such as Flash-Attention [36], Liger-Kernel [37], and gradient checkpointing (activation checkpointing) [15]. Flash-Attention efficiently computes attention without fully materializing the attention matrix, ensuring that peak memory scales linearly rather than quadratically with context length. Liger-Kernel optimizes large intermediate tensor usage during cross-entropy calculation by employing a FusedLinearCrossEntropy mechanism. Notably, the intermediate tensor usage also scales with context length and vocabulary size. Activation checkpointing reduces peak memory by storing only a subset of activations during the forward pass and recomputing them during the backward pass. As context length grows, the volume of checkpointed activations increases, necessitating offload to system memory and on-demand retrieval [15].
Figure 1: Example of long-context CPU offloading with activation checkpointing with a transformer model composed of 4 transformer blocks. Arrows indicate data transfers over PCIe: $P_{i}$ represent model parameters (e.g., attention projection parameters, feed-forward network parameters) for a specific block. $A_{i}$ represents checkpointed input activations for the block. $G_{i}$ represents gradients corresponding to the parameters of the block. The numbered steps illustrate the data movement and computation flow.
Figure 1 illustrates the aforementioned integrated approach, referred to as long-context CPU offloading with offloaded activation checkpointing, using a transformer model with four transformer blocks as an example. Each transformer block contains parameters for attention projections and feedforward layers. The workflow operates as follows: (1) First, necessary parameters are loaded from CPU to GPU memory on a tensor-by-tensor basis. (2) Next, the GPU performs forward computations using these parameters. (3) Checkpointed activations for each transformer block are offloaded to CPU memory. (4) Once the forward pass concludes, the backward pass requires parameters and previously checkpointed activations. (5) These data are then reloaded onto the GPU, which recomputes necessary activations to perform backpropagation. (6) Gradients computed on the GPU are subsequently offloaded to CPU memory, (7) enabling optimizer updates (e.g., using Adam) to execute entirely on the CPU after completing the backward pass. During optimizer steps, full precision parameters, optimizer states, and gradients reside primarily in CPU memory. Such an approach minimizes the volume of data transferred between the GPU and the CPU during each training iteration. Notably, the aforementioned workflow, which combines ZeRo-offload, Flash-Attention, Kiger-Kernel, and offloaded activation checkpointing, is considered the baseline of this study.
To clarify the memory footprint of the workflow shown in Figure 1, the GPU memory usage is first examined. During CPU offloading, the GPU is dedicated solely to computation and retains minimal data: model parameters are streamed block by block, and the corresponding activations and gradients are kept only until each block’s computation is complete. These are then immediately offloaded to system memory or discarded, shifting the primary memory burden to the system memory. On the other hand, the system memory usage is detailed in Table I. The upper half lists components frequently transferred between CPU and GPU during forward and backward passes, while the lower half lists components stored on the CPU for optimizer updates. The memory usage for model parameters and gradients depends on their precision: bf16 requires 2 bytes per parameter, while fp32 requires 4 bytes per parameter. Notably, Zero-Offload uses bf16 on GPUs to manage the huge memory footprint of activations and maximize throughput, while strategically using fp32 on CPU for the sensitive optimizer calculations to ensure the model learns correctly and stably. For checkpointed activations, each GPU requires a unique set of activations; thus, the total system memory usage for these activations is scaled by $N_{g}$. Checkpoints are saved for each transformer block’s input, totaling $L$ blocks, with each activation sized at $B\times C\times H$ elements, stored in bf16 (2 bytes per element). For the Adam optimizer, the optimizer states (momentum and variance) require $8\times P$ bytes in fp32, doubling the memory of gradients due to maintaining two states per parameter. Despite the aforementioned workflow having substantially reduced GPU memory usage, the workflow remains insufficient for long-context fine-tuning. This is because the memory required for activations grows enormously with sequence length. As a result, with parameters, gradients, optimizer states, and these massive activations all resident in system memory, memory pressure escalates rapidly, and system memory itself becomes the dominant bottleneck.
TABLE I: Breakdown of system memory components during CPU offloading. The upper half depicts GPU-CPU transfer size, and the Lower half depicts system memory usage. $P$: total parameters; $N_{g}$: number of GPUs; $B$: batch size per GPU; $C$: context length; $L$: number of transformer blocks; $H$: hidden size.
Component Precision Memory Usage (bytes)
Model parameters bf16 $2\times P$
Gradients bf16 $2\times P$
Checkpointed activations bf16 $2\times(N_{g}\cdot B\cdot C\cdot L\cdot H)$
Model parameters fp32 $4\times P$
Gradients fp32 $4\times P$
Optimizer states fp32 $8\times P$
Figure 2: System memory requirement scaling for 12B across varying context lengths with a batch size of 5.
Figure 3: Throughput and system memory requirement scaling for 12B across batch sizes with a 4K context length.
Figure 4: Comparison of memory access data paths and latencies between local memory and CXL-attached memory
II-B Memory Bottleneck under Long-Context Offloading
In the demanding scenario of training long-context LLMs with CPU offloading techniques, memory demand shifts predominantly from GPU memory to system memory. As a result, system memory capacity becomes a key factor, directly determining the feasible model size, maximum context length, and the batch sizes required to achieve optimal performance. To illustrate this behavior, a motivational experiment was conducted using the 12B model to measure memory requirements and throughput across different context lengths and batch sizes in a 2-GPU setting (hardware specifications are listed in Table II). In the first part of the experiment, the batch size is fixed at 5, and the context length is varied from 512 to 32K tokens. The choice of 32K is based on previous long-context fine-tuning studies [9, 10, 11, 12, 13], which commonly use datasets with context lengths around 32K. For example, LongAlpaca [9] ranges from 3K to 9K, FILM [10] spans 4K to 32K, LongWriter [12] ranges from 2K to 32K, and LongAlign [13] ranges from 8K to 64K, with 90% of samples below 32K. In the second part, the context length is fixed at 4K, while the batch size is varied from 1 to 48 to observe changes in throughput and memory usage. The results are presented in Figures 4 and 4.
Figure 4 shows that CPU memory usage increases linearly with context length. This is because, in long-context CPU offloading, system memory needs to hold checkpointed activations, whose sizes scale proportionally with both the context length and the number of GPUs. Meanwhile, Figure 4 demonstrates that throughput improves with increasing batch size until saturation is reached. This suggests that once the model and context length are fixed, increasing the batch size can enhance GPU utilization. However, Figure 4 shows that CPU memory usage also increases linearly with batch size. This indicates that memory demand is driven not only by model scale and context length but also by batch size when aiming for optimal performance. These findings highlight that in long-context CPU offloading scenarios, system memory usage increases and is likely to become a critical bottleneck as context lengths continue to grow.
Figure 5: Latency of the CPU-based Adam optimizer step with a growing number of parameters.
Figure 6: Comparison of GPU transfer bandwidth from system DRAM and CXL Memory.
(a) Single CXL AIC
(b) Dual CXL AICs
Figure 7: Latency breakdown of CPU offloading: local DRAM baseline vs. naive CXL configurations.
II-C CXL-Attached Memory
Compute Express Link (CXL) is built on top of PCIe and is designed to provide high-bandwidth, low-latency communication between the CPU (host) and various types of devices such as accelerators, memory expanders, and smart I/O devices. CXL differentiates devices into 3 types. Type 1 devices include accelerators with internal caches capable of directly caching host memory. Type 2 devices, such as GPUs, support mutual memory caching with the host, enabling unified memory access. Type 3 devices are intended for memory expansion and include components such as CXL-attached memory for expanding system memory capacity [22]. To enable the use cases of the above devices, three protocol sublayers, including CXL.io, CXL.cache, and CXL.mem, can be combined depending on the device type. While CXL.io provides traditional PCIe-like functionality for configuration, interrupts, and basic I/O operations, CXL.cache and CXL.mem enable devices to transparently cache host memory and allow the host to access memory attached to CXL devices, respectively.
For CXL-attached memory, Figure 4 illustrates the differences in data paths and latency between local and Type 3 devices. Accessing local memory follows a direct path from the CPU cores through the CPU cache and memory controller, resulting in latencies between 80 and 140 nanoseconds [23]. In contrast, accessing CXL-attached memory involves traversing the PCIe interface using the CXL.io and CXL.mem protocols. This path requires coordination between the CPU and the CXL memory controllers, leading to increased latency ranging from 170 to 250 nanoseconds [23]. While CXL-attached memory is utilized as a system memory extension, the Linux kernel integrates CXL-attached memory as CPU-less Non-Uniform Memory Access (NUMA) nodes [33]. This integration allows CXL-attached memory to be managed alongside traditional DRAM, while still enabling users to control allocation explicitly, such as through numactl or libnuma [35], to direct specific data to DRAM, CXL memory, or an interleaved round-robin fashion among available NUMA nodes [35]. Although CXL integration expands usable system memory, our analysis shows that CPU-offloaded long-context LLM fine-tuning remains suboptimal when backed by CXL memory under current deep-learning frameworks (see Section III).
III Observation and Analysis
This chapter empirically characterizes the performance impact of storing CPU-offloaded data on CXL-attached memory. It begins by identifying a critical limitation in current deep learning frameworks that prevents fine-grained memory control. Subsequently, it analyzes the performance of distinct workload components, such as CPU-based computations and GPU data transfers, when using CXL memory. The analysis reveals that naive CXL integration leads to significant end-to-end performance degradation. These findings collectively establish the motivation for the CXL-aware data placement strategies detailed later in this study.
III-A PyTorch Memory Allocation Limitation
PyTorch employs a layered execution stack that lowers high-level Python operations to efficient C++ back-end routines responsible for memory allocation and data movement. This architecture works well in homogeneous memory systems; however, it exposes a central limitation in heterogeneous settings with CXL-attached memory. That is, the current memory placement is controlled at the process level, not at the granularity of individual tensors. In current deployments, NUMA policies configured externally (e.g., via numactl) are applied uniformly to the entire Python process. For example, launching a script with numactl --interleave=0,1 causes all tensor allocations in that process to be interleaved across NUMA nodes 0 and 1, regardless of their role or access pattern. Because PyTorch provides no native mechanism to annotate tensors with placement hints or to route specific allocations to distinct memory tiers, users cannot express policies such as “pin latency-critical optimizer states in local DRAM” while “spilling bandwidth-tolerant or cold tensors (e.g., checkpointed activations) to higher-capacity CXL memory.”
This mismatch is particularly limiting for CPU-offloading pipelines, where different tensor classes exhibit markedly different sensitivity to latency and bandwidth. A single, process-wide policy forces heterogeneous data into a uniform treatment, obscuring opportunities to reduce stall time on the critical path while leveraging capacity elsewhere. The result is suboptimal end-to-end performance and wasted hardware flexibility in mixed-memory environments. These observations motivate a fine-grained allocation extension that (i) exposes per-tensor placement control, (ii) integrates with existing PyTorch allocators without disrupting user code, and (iii) enables principled policies that align tensor characteristics with the most suitable memory tier.
III-B CPU Offloading Slowdowns on CXL-Attached Memory
As illustrated in Section II-A, the CPU‑offloading workflow stores full‑precision parameters, gradients, and optimizer states in system memory so that the CPU can perform the optimizer update locally. As each parameter update is independent, the optimizer phase exhibits ample parallelism. Practical implementations, such as ZeRO-Offload, exploit OpenMP threads and SIMD instructions (e.g., AVX2) to accelerate this compute-intensive step [16]. Consequently, the optimizer phase is highly parallelized and sensitive to increased latency when accessing offloaded data structures. To quantify how memory placement affects optimizer latency within the long‑context CPU‑offloading workflow, the CPU‑based Adam optimizer is benchmarked with offloaded data structures residing either in local DRAM or in CXL-attached memory. Figure 7 summarizes the results. For each configuration, the data size is varied to emulate different LLM scales. Hardware details are listed in Table II, Config. A.
In Figure 7, an “element” consists of a 4-byte parameter, a 4-byte gradient, and 8 bytes of optimizer state. The SIMD kernel processes each element in three steps: (i) it loads the parameter, gradient, and state from memory into vector registers; (ii) executes the floating‑point update; and (iii) writes the updated values back. For modest data volumes, the latency penalty of CXL‑attached memory is negligible; however, once the element count exceeds roughly 20 million, optimizer time with CXL-attached memory rises sharply, reaching nearly 4 times the DRAM baseline. The primary cause is the higher access latency of the CXL path (170-250 ns) compared to local DRAM (80-140 ns), as shown earlier in Figure 4. These results indicate that naively placing latency‑critical optimizer data in CXL‑attached memory can severely degrade fine‑tuning performance. Effective CXL deployments for long‑context CPU‑offloading need to keep latency‑sensitive data in low‑latency DRAM and relegate latency‑tolerant data to CXL‑attached memory, respectively.
III-C GPU Data Transfers on CXL-Attached Memory
The CPU-offloading workflow involves not only a CPU-intensive optimizer step but also frequent, high-volume data transfers between system and GPU memory. For example, before each layer’s backward computation, model parameters and checkpointed activations are copied from system memory to the GPU. Afterward, the resulting gradients are copied back. To quantify the impact of physical memory location on data transfer behavior, this study evaluates GPU copy performance with source buffers placed in either local DRAM or CXL-attached memory. In each experiment, page-aligned host buffers are allocated and pinned to a specific NUMA node. These buffers are registered to enable direct DMA transfers over PCIe, thereby bypassing intermediate copies through CPU caches. Asynchronous memory transfers are then issued, and the resulting effective bandwidth is measured. Figure 7 shows the results.
Figure 8: System architecture of the proposed methods, in which a fine-grained allocation extension adds per-tensor placement control and a CXL-aware memory allocator directs tensors to local DRAM or CXL-attached memory within the existing software stack.
As shown in the figure, the observed transfer bandwidth from CXL-attached memory closely matches that of local DRAM. This is attributed to the use of direct DMA over PCIe, which allows CXL memory to bypass the CPU and transfer directly to the GPU. Since local DRAM also relies on PCIe for such transfers, both memory types share a similar topology. Throughput increases with transfer size until it saturates the PCIe interface bandwidth. This convergence occurs because page-locked buffers expose equivalent DMA paths across both memory types, making the operation bound by interface limits. Furthermore, GPU data transfers tend to tolerate latency more effectively. This is because, for example, CPU-offloading systems can utilize prefetching or asynchronous offloading to mask latency. These results indicate that for high-throughput, latency-tolerant operations such as GPU transfers, CXL-attached memory can deliver performance comparable to local DRAM. This contrasts with latency-sensitive phases, such as optimizer steps, which reinforce the importance of workload-aware data placement.
III-D End-to-End Fine-Tuning Slowdown
Based on the characteristics outlined in previous sections, this section measures how naively incorporating CXL-attached memory affects end-to-end performance during LLM fine-tuning. The comparison evaluates a local DRAM baseline against configurations that combine local DRAM with CXL memory under a naive interleaving policy. This naive approach is representative of what occurs due to the PyTorch memory allocation limitations described in Section III-A. Figure 7 presents the latency profile for fine-tuning a 12-billion-parameter model. In Figure 7(a), a single CXL AIC is used. The optimizer step, labeled as the STEP phase, suffers the most significant slowdown. This is because its CPU-bound loads and stores are acutely sensitive to the higher access latency of CXL memory, a direct consequence of the naive interleaving policy placing critical optimizer data on the slower tier. Phases dominated by GPU transfers, specifically FWD and BWD, exhibit smaller slowdowns because prefetching and asynchronous DMA obscure some of the added latency.
Figure 7(b) shows the result of adding a second CXL AIC. With more available bandwidth from the additional device, the overall performance improves, and the slowdown is mitigated compared to the single-AIC case. However, a performance gap still remains relative to the DRAM-only baseline. This demonstrates that while adding more hardware can alleviate bandwidth issues, it does not resolve the fundamental problem of latency sensitivity in the optimizer step. The naive, process-level memory policy remains a bottleneck. These findings underscore the need for a more intelligent, CXL-aware data placement strategy. Such a strategy needs to distinguish between latency-sensitive and latency-tolerant data and leverage the aggregate bandwidth of multiple CXL AICs effectively.
IV CXL-Aware Long-Context LLM Fine-Tuning
To address the performance degradation arising from the naive adoption of CXL-attached memory, this work introduces a two-part methodology that overcomes the limitations of existing deep learning frameworks and intelligently manages data placement in heterogeneous memory systems. The first component is a fine-grained, per-tensor memory allocation extension for PyTorch, providing essential control over NUMA memory policies. The second is a CXL-aware memory allocator that leverages this extension to strategically place tensors based on their latency sensitivity. Figure 8 illustrates the integration of these components into the LLM fine-tuning stack, where the extension operates between Python and C++ layers of PyTorch, and the CXL-aware allocator directs memory decisions from a higher level.
IV-A Fine-grained Memory Allocation Extension
As discussed previously, on the path to exploiting CXL-attached memory under the CPU-offloading scenario, a significant limitation of existing deep learning frameworks, such as PyTorch, lies in their coarse-grained memory management. By default, NUMA policies are enforced at the process level, causing all tensors allocated within a script to share the same placement strategy. This design prevents applications from selectively assigning tensors to different memory tiers according to their individual access patterns and latency sensitivities. To address this limitation, a fine-grained memory allocation extension for PyTorch is developed to introduce per-tensor control over NUMA placement, allowing developers to allocate tensors directly to local DRAM, a designated CXL device, or an interleaved set of nodes. The implementation is shown in Figure 9. It consists of a Python wrapper that interfaces with C++ backend functions and uses the libnuma library, specifically, numa_alloc_onnode for single-node placement and numa_alloc_interleaved_subset for customized interleaving across nodes. To support efficient GPU data transfers, the extension leverages cudaHostRegister from the CUDA toolkit to pin host memory, thereby enabling zero-copy direct memory access (DMA), while torch.from_blob is used to construct a PyTorch tensor that references the externally managed memory block without additional data copies. This extension establishes the foundational mechanism required for the CXL-aware memory allocator and enables future exploration of advanced memory management strategies in heterogeneous systems.
Figure 9: Workflow of the fine-grained memory allocation extension for PyTorch.
IV-B CXL-Aware Memory Allocator
To effectively exploit the aforementioned fine-grained, per-tensor control for the PyTorch memory allocation extension, this study further develops a CXL-aware memory allocator, which is a runtime algorithm that dynamically partitions CPU-offloaded data between local DRAM and CXL-attached memory. The allocator is designed to minimize performance degradation caused by CXL’s higher access latency by prioritizing data placement based on each component’s latency sensitivity. Its design builds on two key ideas: (1) hierarchical grouping of data by latency tolerance, (2) a latency-first greedy allocation strategy that enforces this hierarchy. In Section IV-B3, the adaptive placement behavior is demonstrated under varying hardware configurations and memory pressure to demonstrate its effectiveness. Together, these mechanisms enable efficient and transparent utilization of heterogeneous memory resources during long-context LLM fine-tuning.
IV-B1 Latency Sensitivity Classification
The allocator classifies CPU-resident data into latency-tolerance levels based on access patterns observed during fine-tuning. Comparing CPU and GPU access behaviors shows that data both accessed and computed on the CPU lies on the critical path and is therefore highly latency-sensitive. For GPU accesses under CPU-offloaded training, this study observes two behaviors: fetch (CPU→GPU transfers required before a layer’s computation) and offload (GPU→CPU transfers performed after computation completes). Because training proceeds layer by layer, fetches must complete immediately (e.g., prefetching bf16 weights for the next forward or backward layer), whereas offloads can be asynchronous (e.g., checkpointed activations after the forward pass, or gradients after the backward pass). Based on the fetch and offload behaviors, three conditions are defined, which are $C_{1}$ compute on the CPU, $C_{2}$ fetching by the GPU, and $C_{3}$ offloading by the GPU. These conditions induce four latency-tolerance levels used by the allocator.
  • Level 1 (Lowest tolerance). Allocations satisfying $C_{1}$ (tight load–compute–store loops on the CPU), such as optimizer states, master weights, and master gradients.
  • Level 2 (Low tolerance). Allocations satisfying only $C_{2}$ (GPU fetch), for example, bf16 weights prefetched prior to a layer’s forward or backward pass.
  • Level 3 (Medium tolerance). Allocations satisfying both $C_{2}$ and $C_{3}$, where accesses are partly hidden by prefetch and asynchronous offload; for example, checkpointed activations offloaded after the forward pass and later fetched for recomputation during backpropagation.
  • Level 4 (Highest tolerance). Allocations satisfying only $C_{3}$ (purely asynchronous offload) that do not lie on the immediate critical path, such as per-layer gradients transferred to CPU memory after computation.
Input: $S_{\text{local}}$: Local DRAM size;
$S_{\text{cxl}}$: Aggregated CXL size;
$num_{\text{cxl}}$: Number of CXL devices;
group_items_sizes: maps each latency level to its list of item sizes $\{level:[size_{1},\dots]\}$
Output: allocations: item $\rightarrow$ (policy, interleave ratio)
$S_{\text{remain}}\leftarrow S_{\text{local}}$;
allocations $\leftarrow\{\}$;
for $level\leftarrow 1$ to 44 do
 foreach $size_{i}$ in group_items_sizes[level]group\_items\_sizes[level] do
    $key\leftarrow$ "level_levellevel_item_ii";
    if $S_{\text{remain}}\geq size_{i}$ then
       allocations[keykey] $\leftarrow$ (pure_local, 1:01{:}0)
       $S_{\text{remain}}\leftarrow S_{\text{remain}}-size_{i}$;
    else if $S_{\text{remain}}>0$ then
       allocations[keykey] $\leftarrow$ (local_cxl, $(1{:}\underbrace{1{:}\dots:1}_{num_{\text{cxl}}})$);
       $S_{\text{remain}}\leftarrow 0$;
    else
       allocations[keykey] $\leftarrow$ (pure_cxl, $\underbrace{1{:}\dots:1}_{num_{\text{cxl}}}$);
return allocations;
Algorithm 1 Latency-First Allocation
Figure 10: Allocation example for single CXL device and sufficient DRAM.
Figure 11: Allocation example for multiple CXL devices and sufficient DRAM.
Figure 12: Allocation example for multiple CXL devices and limited DRAM.
IV-B2 Latency-First Greedy Algorithm
Building upon the four-level latency classification, the latency-first greedy algorithm determines how CPU-resident data are distributed across DRAM and multiple CXL-attached memory devices. The guiding principle of the algorithm is to prioritize low-latency data placement in local DRAM while progressively offloading or interleaving less latency-insensitive data into CXL memory as DRAM capacity becomes limited. This design follows a greedy strategy, which makes locally optimal decisions at each step, so that latency-critical data always receive preferential treatment without incurring the complexity of global optimization or iterative tuning. Algorithm 1 presents the detailed allocation process. The algorithm takes as input the available local DRAM capacity ($S_{local}$), the aggregated capacity of all CXL devices ($S_{cxl}$), and the number of CXL devices ($num_{cxl}$). The CPU-resident data are grouped by latency sensitivity into the mapping structure group_items_sizes, which associates each of the four tolerance levels with a list of tensor or parameter sizes. The output is an allocation table that maps each data item to a specific placement policy and its corresponding interleave ratio across memory tiers.
The allocator begins with the full DRAM capacity ($S_{remain}$$S_{local}$) and iterates through the latency levels from the most sensitive classification (Level 1) to the least sensitive classification (Level 4). Within each level, every item is evaluated based on its size and the remaining DRAM space. If an item completely fits within the available DRAM ($S_{remain}\geq Size_{i}$), it is placed entirely in local DRAM, marked as pure_local, ensuring minimal latency for the most performance-critical components such as optimizer states and master gradients. If DRAM space is insufficient but non-zero ($S_{remain}>0$), the allocator interleaves the data across the remaining DRAM and all CXL devices, adopting the local_cxl policy with an even interleave ratio. This configuration allows the DRAM portion to serve as a fast buffer while using CXL memory to absorb the overflow and take advantage of the bandwidth from multiple CXL memory devices. Once DRAM is exhausted ($S_{remain}=0$), the remaining items are distributed evenly across all CXL devices using the pure_cxl policy, thereby maximizing aggregate bandwidth and overall capacity.
This greedy traversal ensures that DRAM is always reserved for the most latency-critical tensors while maintaining balanced utilization of the heterogeneous memory system. The resulting allocation map explicitly specifies both the placement policy and interleaving configuration, allowing the runtime system to reproduce consistent memory behavior across runs. By following a latency-first decision order, the allocator achieves an effective compromise between minimizing response time for critical data accesses and exploiting the extended capacity of CXL memory. Furthermore, its deterministic, lightweight design makes it suitable for integration into existing deep-learning runtimes without requiring online profiling or costly dynamic migration during fine-tuning.
IV-B3 Allocation Examples
The behavior of the CXL-aware allocator is illustrated through three representative scenarios that reveal how the algorithm adapts to different hardware configurations and memory pressures.
  • Scenario 1 (Single CXL Device, Ample DRAM). As shown in Figure 12, a system equipped with one CXL device and sufficient DRAM places all latency-critical Level 1 data, which includes optimizer states (fp32 O), master parameters (fp32 P), and master gradients (fp32 G), entirely in DRAM. The more latency-tolerant data, such as checkpointed activations (bf16 A), parameters (bf16 P), and gradients (bf16 G), are stored in the CXL memory.
  • Scenario 2 (Multiple CXL Devices, Sufficient DRAM). In Figure 12, multiple CXL devices are introduced while DRAM capacity remains sufficient for Level 1 data. Latency-sensitive data stay in DRAM, whereas latency-tolerant data (bf16 A, bf16 P, bf16 G) are interleaved across the CXL devices to exploit their aggregate bandwidth, enhancing throughput for GPU data transfers.
  • Scenario 3 (Multiple CXL Devices, Limited DRAM). As illustrated in Figure 12, when DRAM is insufficient to accommodate all Level 1 data, the greedy allocator first fills the available DRAM with the highest-priority tensors. The remaining Level 1 data, such as part of the optimizer states (fp32 O), are striped across the remaining DRAM and all CXL devices. This fallback mechanism ensures balanced utilization of memory resources while minimizing the performance penalties associated with higher-latency CXL access.
(a) 7B model in a single-GPU scenario
(b) 7B model in a dual-GPU scenario
(c) 12B models in a single-GPU scenario
(d) 12B models in a dual-GPU scenario
Figure 13: Training throughput comparison of three single-AIC configurations: Baseline (local DRAM only), Naive CXL (interleaving), and CXL-Aware Allocation (labeled as Our), evaluated across varying models, context lengths and batch sizes.
V Evaluations
V-A Experimental Setup
V-A1 Hardware and Software Specification
Experiments run on a server whose hardware and software stack are summarized in Table II. The platform combines a high‑performance CPU (e.g., Intel Xeon 6780E), ample local DRAM (e.g., 512 GB DDR5), and two cutting‑edge GPUs (e.g., NVIDIA H100) optimized for LLM workloads. CXL memory expansion is evaluated in two configurations: a single‑AIC setup and a dual‑AIC setup designed to reveal scalability limits and bandwidth‑contention effects. Both add‑in cards were SMART Modular CMM-CXL-2.0 devices [25]. Notably, even though the current configuration is limited to the availability of CXL AICs, we argue that the observations made in this study are representative and transferable, as they are governed by the fundamental bandwidth and latency characteristics of the CXL protocol itself.
V-A2 Workload Setup
Data placement across local DRAM and CXL‑attached memory is managed by libnuma, which is exposed to PyTorch through a lightweight custom extension that intercepts memory allocation calls to enforce NUMA policies. DeepSpeed [38], which is the implementation of ZeRO-Offload, handles CPU offloading, while Flash‑Attention, Liger‑Kernel, and activation checkpointing enable efficient long‑context processing. Checkpointed activations, once generated, are offloaded to host DRAM. The study focuses on fine‑tuning large language models under a Causal Language Modeling objective. Two representative models, including Qwen2.5‑7B [39] and Mistral NeMo 12B [40], serve as workloads for exploring performance and scalability across varying context lengths, batch sizes, GPU counts, and AIC configurations. All experiments employ bf16 mixed‑precision training. The Adam optimizer maintains fp32 master parameters and optimizer states on the CPU, delivering the numerical stability required for LLM fine‑tuning. Details on specific context lengths, batch sizes, and AIC setups appear in the individual results sections.
TABLE II: Hardware and Software Specification for Experimental Setup.
Component Specification
Hardware
OS Ubuntu 24.04 LTS
Linux Kernel v6.9
CPU 1 $\times$ Intel(R) Xeon(R) 6780E
GPUs 2 $\times$ NVIDIA H100 80GB PCIe
PCIe PCIe 5.0 (x16 links for GPUs and AICs)
Local DRAM 512 GB (4 $\times$ 128 GB DDR5-6400)
CXL AICs. (Config. A) 1 $\times$ CXA-8F2W (512 GB AIC)
CXL AICs. (Config. B) 2 $\times$ CXA-4F1W (256 GB AIC)
Software
NUMA Control numactl, libnuma 2.0.19
PyTorch torch 2.5.1
Model transformers 4.47.1
Framework deepspeed 0.16.2
(a) 7B model in a single-GPU scenario
(b) 7B model in a dual-GPU scenario
(c) 12B models in a single-GPU scenario
(d) 12B models in a dual-GPU scenario
Figure 14: Training throughput comparison of three dual-AIC configurations: Baseline (local DRAM only), Naive CXL (interleaving), and CXL-Aware Allocation (labeled as Our), evaluated across varying models, context lengths and batch sizes.
V-B Performance Evaluation with a Single CXL AIC
The single-AIC setup, corresponding to Config. A in Table II, serves as the baseline environment. Three configurations are evaluated to quantify the impact of CXL-attached memory. The first is the baseline, where all data reside in local DRAM. The second, naive CXL, combines 128 GiB of local DRAM with 512 GiB of CXL memory using a uniform numactl --interleave=all policy. The third, CXL-aware allocation, employs the same capacities but applies the proposed fine-grained memory allocation extension and CXL-aware memory allocator.
Figure 13(a) shows single-GPU throughput for a 7 B model under varying context lengths (4 K–32 K) and batch sizes (1–32). Relative to the baseline (normalized to 100%), the naive CXL configuration sustains only 76%–94% throughput, depending on workload mix. Workloads dominated by forward (FWD) and backward (BWD) passes experience smaller losses because the latency-sensitive optimizer (STEP) occupies a smaller runtime fraction. By contrast, CXL-aware allocation restores throughput to 97%–99%, reducing the degradation to just 1%–3% compared to DRAM-only and outperforming the naive policy by up to 21%.
Figure 13(b) presents dual-GPU results for the same 7 B model. The naive CXL setup achieves 77%–93% of baseline throughput, while CXL-aware allocation improves this to 93%–97%, narrowing the gap to 3%–7%. The smaller gain compared with the single-GPU case stems from bandwidth contention: both GPUs share a single AIC, effectively halving available read/write bandwidth. This bandwidth constraint limits both FWD and BWD phases, underscoring the benefit of deploying multiple AICs. Even with identical total capacity, multiple AICs aggregate bandwidth more effectively, allowing each GPU to access sufficient bandwidth and thus recover full performance. Section V-C further analyzes this multi-AIC configuration.
Figure 13(c) extends the analysis to a 12 B model under a single-GPU setup. The naive CXL configuration reaches 72%–93% of baseline throughput, while CXL-aware allocation raises this to 88%–96%, reflecting a 4%–12% shortfall relative to DRAM-only but up to 16% improvement over naive CXL. Although the benefit remains evident, the smaller margin arises from latency sensitivity rather than bandwidth limits. In this case, local DRAM cannot hold all latency-critical data, forcing some onto the slower AIC. This co-location increases access latency, which is alleviated when dual AICs are employed—providing additional bandwidth that effectively reduces latency. Section V-C presents detailed results for this case, where CXL-aware allocation can even surpass the DRAM-only baseline.
Finally, Figure 13(d) reports dual-GPU throughput for the 12 B model with one AIC. The naive CXL setup sustains 72%–91% of baseline performance, while CXL-aware allocation increases it to 82%–92%, offering up to 10% improvement over the naive policy. Nonetheless, concurrent bandwidth contention and latency sensitivity constrain further gains. The following section evaluates how the proposed allocation algorithm alleviates these bottlenecks under a dual-AIC environment.
V-C Performance Evaluation with Dual CXL AICs
The evaluation next turns to the dual-AIC scenario, designated as Config. B in Table II. Three configurations are used to establish the impact of CXL memory. The first is the baseline, where all data remains in local DRAM. The second is naive CXL, which pairs 128 GiB of DRAM with two 256 GiB AICs under a naive numactl --interleave=all policy. The third is CXL-aware allocation, which uses the same capacities but applies the proposed algorithm.
Figure 14(a) presents single-GPU throughput for a 7 B model across the same range of context lengths and batch sizes used earlier in the single-AIC configuration. While the naive CXL policy results in a 2% to 9% performance drop compared to the baseline, the proposed CXL-aware allocation restores performance to 100% of the baseline, showing that no performance is lost when CXL memory is managed intelligently. The algorithm achieves this by automatically interleaving latency-tolerant data across both AICs, thereby capitalizing on their aggregate bandwidth. These results demonstrate that with workload-aware allocation, a dual-card CXL configuration can fully match native DRAM performance.
Figure 14(b) shows dual-GPU throughput for a 7 B model across the same context lengths and batch sizes used in the single-AIC setup. The naive CXL policy lowers performance by 2% to 8% relative to the baseline, whereas the CXL-aware allocation restricts the loss to no more than 1%. In the single-AIC case shown in Figure 13(b), bandwidth contention remained a limiting factor even with CXL-aware placement. The dual-AIC setup alleviates this issue, allowing the algorithm to exploit the combined resources of both cards and largely eliminate the penalty, while the naive policy continues to falter due to poor placement.
Figure 14(c) illustrates single-GPU throughput for a 12 B model across the same range of context lengths and batch sizes. The naive CXL policy reduces throughput by 2% to 11% compared to the baseline, whereas the CXL-aware allocation restores performance to 100%–101% of the baseline. In the single-AIC case (Figure 13(c)), limited local DRAM capacity constrained performance even with CXL-aware placement. That limitation still exists in the dual-AIC setup, but the CXL-aware allocation mitigates its impact by leveraging the combined bandwidth of both cards alongside local DRAM. The latency-first greedy algorithm places latency-sensitive data in local DRAM whenever possible and assigns the remaining data across local DRAM and both AICs to aggregate bandwidth and minimize overall memory-access latency. The naive CXL policy, by contrast, continues to underperform due to unoptimized placement.
Finally, Figure 14(d) reports dual-GPU throughput for the 12 B model. The naive CXL policy reduces throughput by 2% to 9% relative to the baseline, while the CXL-aware allocation limits the loss to at most 1%. In the single-AIC case (Figure 13(d)), bandwidth contention and limited DRAM capacity remained bottlenecks even with CXL-aware placement. The dual-AIC configuration resolves these issues: CXL-aware allocation places latency-sensitive data appropriately and distributes remaining allocations across all CXL NUMA nodes to fully aggregate bandwidth. This automated interleaving enables both CPU and GPU workloads to exploit the combined bandwidth of the dual cards. The resulting performance, effectively matching the DRAM-only baseline, underscores the importance of this work to intelligently orchestrate underlying hardware resources.
(a) 7B model in a single-GPU scenario
(b) 7B model in a dual-GPU scenario
(c) 12B models in a single-GPU scenario
(d) 12B models in a dual-GPU scenario
Figure 15: Training latency decomposition for a long-context workload (32K context, batch size 4), showing our CXL-aware allocation consistently outperforms the naive CXL policy and nearly matches the DRAM-only baseline performance on both single- and dual-AIC systems. Detailed training latency comparison for five configurations, sorted by performance from right to left. The configurations are: (1) baseline, using local DRAM only, (2) naive CXL with single-AIC, (3) CXL-aware allocation with a single-AIC, (4) naive CXL with dual-AIC, and (5) CXL-aware allocation with dual-AIC.
V-D Latency Decomposition
While the previous section established end-to-end performance, this section breaks down the latency of FWD, BWD, and STEP for each configuration to further analyze latency and the impact of CXL-aware allocation. The system setup and configurations follow those used earlier: the single-AIC scenario corresponds to Config. A in Table II, and the dual-AIC scenario corresponds to Config. B. Five configurations are compared. The first is the baseline, where all data remain in local DRAM. The second is naive CXL (1 AIC), which combines 128 GiB of local DRAM with 512 GiB of CXL memory under a naive numactl --interleave=all policy. The third is CXL-aware allocation (1 AIC), which uses the same capacities but applies the proposed CXL-aware memory allocator. The fourth is naive CXL (2 AIC), pairing 128 GiB of DRAM with two 256 GiB AICs under the same interleave-all policy. The fifth is CXL-aware allocation (2 AIC), which employs the same capacities while applying the proposed extension and allocator.
Figure 15 presents a detailed latency breakdown of the training process into forward (FWD), backward (BWD), and optimizer-step (STEP) phases, revealing how memory configurations affect each stage. In the 7 B single-GPU scenario shown in Figure 15(a), FWD and BWD remain around $8.5s$ and $29.3s$ across all policies. The major separation arises in STEP: the baseline completes in $\approx 1.95s$, naive CXL (1 AIC) inflates it to $\approx 4.06s$, and naive CXL (2 AIC) reduces it slightly to $\approx 2.58s$. In contrast, CXL-aware allocation lowers STEP to $\approx 2.08s$ on one AIC and $\approx 1.96s$ on two AICs, essentially matching the baseline. This is because the naive interleave policy places latency-critical optimizer states on AIC memory, forcing frequent CPU accesses through a high-latency path. On the other hand, the CXL-aware allocation pins latency-sensitive tensors in local DRAM and pushes only bandwidth-tolerant data to CXL, eliminating this penalty.
Figure 15(b) shows the 7 B dual-GPU scenario, where bandwidth contention on the CXL interconnect becomes more pronounced. With a single AIC, both GPUs compete for bandwidth, slowing FWD and BWD even under CXL-aware allocation (1 AIC). This bottleneck highlights the limitation of a single CXL device in a multi-GPU environment. However, CXL-aware allocation (2 AIC) resolves this issue by intelligently distributing memory accesses across both AICs, providing sufficient aggregate bandwidth and restoring performance to within 0.2% of the DRAM-only baseline. For the larger 12 B model in a single-GPU setup (Figure 15(c)), FWD and BWD remain similar across configurations, but STEP dominates. The baseline STEP is $\approx 2.57s$; naive CXL (1 AIC) raises it to $\approx 6.11s$, while CXL-aware allocation (1 AIC) lowers it to $\approx 3.73s$. CXL-aware allocation (2 AIC) further cuts STEP to $\approx 2.59s$, effectively equal to the baseline. The small residual gap in the single-AIC aware case aligns with earlier findings: local DRAM cannot accommodate all latency-critical tensors for 12 B, so part of STEP still touches AIC memory. With only one card, this additional latency remains unavoidable.
Finally, the most demanding configuration, the 12 B model with dual GPUs, shown in Figure 15(d), demonstrates the combined effects of latency sensitivity and bandwidth contention. Single-AIC configurations suffer substantial degradation, with total latency increasing by up to 9% over the baseline because the AIC is fully saturated. Both FWD and BWD phases, as well as STEP, slow down. In contrast, CXL-aware allocation (2 AIC) effectively manages the hardware resources: by distributing allocations across local DRAM and both CXL NUMA nodes, it aggregates bandwidth while prioritizing latency-sensitive data, eliminating CXL-induced overhead and matching the performance of the DRAM-only baseline.
VI Related Works
Offloading strategies have emerged as an effective means of breaking the GPU memory ceiling, enabling the training of models that would otherwise exceed on-device capacity by staging tensors in CPU DRAM or NVMe SSDs [16, 17, 18, 19, 20]. Among these, the ZeRO series [16, 17] has become the most widely adopted, owing to continuous maintenance that fixes bugs, adds support for new models, and preserves interoperability with complementary optimizations such as Flash-Attention [36] and Liger-Kernel [37]. Its reference implementation now underpins several popular training frameworks, including Accelerate [41] and MS-Swift [42]. Complementary work further broadens the design space of memory-centric training. Huang et al. [18] tailor an offloading mechanism specifically for large-scale language models, while Chen et al. [19] evaluate pragmatic orchestration policies that coordinate CPU and GPU memory during fine-tuning. Zeng et al. [20] extend the idea to fully heterogeneous environments, automatically balancing both compute and memory resources.
For CXL-attached memory, several tiered-memory systems (TMS) address the capacity challenge at the operating-system or hardware level. Representative examples include TPP [32], which classifies pages as hot or cold for placement, and NOMAD [43], which employs transactional page migration for asynchronous data movement. Other advanced systems, such as Colloid [44], dynamically balance traffic to equalize effective latency, while M5 [45] embeds hardware trackers in the CXL controller to provide fine-grained access statistics. Although these general-purpose TMS designs operate transparently without requiring application modifications, their workload-agnostic nature can lead to suboptimal performance for specialized applications such as LLM fine-tuning.
The analysis in this study identifies two weaknesses of applying a generic TMS to this workload. First, for data components transferred to the GPU, CXL-attached memory already offers bandwidth comparable to local DRAM. A TMS that migrates these pages to DRAM before a transfer would incur unnecessary page-movement overhead without improving performance. Second, the optimizer step exhibits a streaming access pattern in which every data element is updated once per iteration. This uniform pattern lacks the distinct hot or cold data regions that TMS architectures are designed to exploit. A generic tiering system would therefore be ineffective and could even introduce additional overhead through misguided migrations. In contrast, this study leverages application-specific knowledge to statically partition data, providing a more effective strategy for the predictable memory-access patterns of fine-tuning workloads.
VII Conclusion
To address the latency challenge of CXL-attached memory relative to local DRAM during long-context LLM fine-tuning, particularly the performance degradation observed in CPU-intensive computations, this study identifies a fundamental limitation in existing deep learning frameworks: the lack of fine-grained control over memory allocation across heterogeneous memory systems. To overcome this limitation, two complementary approaches are proposed. First, a fine-grained memory allocation extension to PyTorch is developed to provide tensor-level control over system memory allocation policies, enabling fine-grained tensor management and facilitating future research with CXL-attached memory. Second, a CXL-aware memory allocator is introduced, which strategically places latency-sensitive data in local DRAM and latency-tolerant data in CXL-attached memory based on their access patterns. Together, these optimizations substantially mitigate the performance drawbacks of CXL-attached memory, consistently outperforming naive CXL adoption with up to 21% improvement across evaluated scenarios. In the dual-AIC configuration, the proposed method achieves near-baseline throughput, narrowing the gap with DRAM-only setups to within 1%, thereby demonstrating that CXL-attached memory can serve as a practical and high-performance solution for long-context LLM fine-tuning.